A partial result of the computation for a two-factor factorial experiment is given as follows (assume that the number of experimental units in each group is fixed):

Source of Variation SS df MS F
Factor A 48 4 12 1
Factor B 6 5
Interaction (AB) 24 15
Error 105 -
Total 139 - -

A) Fill in all the missing results in the ANOVA table.

B) At the 0.05 level of significance, is there an interaction effect? (Show your work: your hypotheses, test statistic, critical value, and decision).

C) At the 0.05 level of significance, is there evidence of an effect due to factor A? (Show your work: your hypotheses, test statistic, critical value, and decision).
Answer #1

Part A

MS A / MS error = F

12/a = 1

a = 12

MS error = 12

SS/df = MS

MS interaction = 15

15 = SS interaction / 24

SS interaction = 15*24 = 360

MS B / MS error = 5

MS B /12 = 5

MS B = 60

Required ANOVA table is given as below:

Source of Variation

SS

df

MS

F

Factor A

48

4

12

1

Factor B

360

6

60

5

Interaction (AB)

360

24

15

1.25

Error

1260

105

12

-

Total

2028

139

-

-

Part b

H0: There is no interaction.

Ha: There is an interaction.

α = 0.05

Test statistic = F = MS interaction / MS error = 60/12 = 5

Critical value = 1.621485

Test statistic > Critical value

So, we reject the null hypothesis

There is sufficient evidence to conclude that there is a statistically significant interaction.

Part c

H0: There is no significant effect due to factor A.

Ha: There is a significant effect due to factor A.

α = 0.05

F = 1

Critical value = 2.45821

Test statistic F is less than critical value

So, we do not reject the null hypothesis

There is not sufficient evidence to conclude that there is significant effect due to factor A.